# Comparing Hubbard parameters from linear-response theory and Hartree-Fock-based approach

### v1

@@ RESULTS
Figure 3 extends our analysis by including a Hubbard correction on the oxygen 2pp states ($U_{O}$), computed with both LRT and eACBN0. Within LRT, the resulting $U_{O}$ values are large ($\sim 10$ eV), consistently with previous studies KirchnerHall:2021. This is a consequence of the high electronegativity of oxygen: its 2pp shell is nearly fully occupied, which leads to a weak electronic response and therefore a large $U$ in the LRT framework Gelin:2024, Yu:2014. Despite these large $U_{O}$ values, the inter-site interaction $V_{TM,O}$ remains small (below 1 eV), indicating that TM–O hybridization has only a minor effect on the electronic structure of these TMOs even when $U_{O}$ is considered. In LRT, $U_{TM}$ and $U_{O}$ appear as diagonal elements, and $V_{TM,O}$ as off-diagonal elements of the matrix obtained by inverting the interacting and non-interacting response matrices, $\chi$ and $\chi_{0}$ [Eqs. (6) and (7)]. Including oxygen among the perturbed subspaces alters the full response matrices, because screening channels involving O are now explicitly taken into account. Once both TM and O sites are perturbed, the structure of the response matrices in DFT+$U$ and DFT+UU+$V$ becomes effectively similar: both schemes capture the same screening pathways involving TM and O states. Consequently, the value of $U_{TM}$ can change when $U_{O}$ is included in the LRT calculation. Comparing Figs. 2 and 3, we observe a slight increase of $U_{TM}$ for CoO and NiO (by 3355%), a substantial increase for FeO (18%), and almost no change for MnO. Conversely, eACBN0 yields smaller $U_{O}$ values than LRT ones, in the range of 232-3 eV. Similar to the case with the $V_{TM,O}$ correction, the converged Hubbard parameters can vary depending on the Hubbard manifolds being considered. In this self-consistent cycle, the parameters from one iteration feed back into the next as a correction term, thereby influencing the KS states and the final converged values. Comparing Figs. 2 and 3 for eACBN0, we find (consistently with LRT) that $U_{TM}$ in CoO increases slightly (by 1122%), while in FeO it slightly decreases (by 1144%) in contrast to LRT. For NiO and MnO, $U_{TM}$ exhibits some variations (small increases, decreases, or no significant change), depending on whether the DFT+$U$ or DFT+UU+$V$ scheme is employed.
@@
Finally, following the protocol in Fig. S1 in the SI Timrov:2021, we compute the self-consistent Hubbard parameters and their corresponding optimized structural parameters using both LRT and eACBN0. The results are presented in Fig. S3 and S4 and their discussion is in the SI. For both methods, the structural parameters (lattice constants and rhombohedral angles) are generally in good agreement with available experimental data.
@@
For the inter-site interactions between Zr-4dd and O-2pp states ($V_{Zr,O}$), LRT yields values below 1 eV, comparable to those obtained for the TMOs discussed in the previous section. Including these interactions in the DFT+UU+$V$ calculation increases the band gap by 7% relative to DFT+$U$ with only $U_{Zr}$ from LRT. In contrast, eACBN0 produces substantially larger $V_{Zr,O}$ values (2.3–2.5 eV), leading to a 22% band-gap increase compared to DFT+$U$ with only $U_{Zr}$, as shown in Fig. 4(b). Since $V_{Zr,O}$ depends on the interatomic distances between Zr and its nearest-neighbor O atoms, a range of values is reported in Fig. 4(a) rather than a single number. Despite these $V_{Zr,O}$ corrections, both methods still underestimate the experimental band gap by more than 1 eV Bersch:2008.
@@
Figure 4: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZrO2 without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Zr}$ parameter for 4dd orbitals of Zr ions and inter-site Hubbard $V_{Zr,O}$ parameter between 4dd orbitals of Zr ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap Bersch:2008 is shown by the dashed blue line in panel (b).
Next, we turn to BaTiO3, which is similar to ZrO2 in the sense that Ti4+ also has a nominal $d^{0}$ configuration. In this case, we consider not only the band gap but also the crystal-structure properties, which are highly sensitive to Hubbard corrections Gebreyesus:2023, Choi:2025. BaTiO3 is a well-known ferroelectric material, exhibiting a rhombohedral distortion at low temperatures. However, standard DFT+$U$ with sizable $U$ corrections for Ti-$3d$ states incorrectly stabilizes a centrosymmetric cubic structure, thereby suppressing its intrinsic polarization Tsunoda2019prm. However, including the inter-site interactions ($V_{Ti,O}$) between Ti-$3d$ and O-$2p$ states preserves their hybridization, which is essential for maintaining the rhombohedral distortion and thus restoring the polarization Gebreyesus:2023, Choi:2025.
@@
Figure 6: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for BaTiO3 without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Ti}$ parameter for 3dd orbitals of Ti ions and inter-site Hubbard $V_{Ti,O}$ parameter between 3dd orbitals of Ti ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap Wemple:1970 is shown by the dashed blue line in panel (b).
@@
When $U_{O}$ is included, the structural behavior changes, as shown in Figs. S5(e)–(h). In this case, LRT fails to capture the ferroelectric nature of BaTiO3: the large $U_{Ti}$ and $U_{O}$ values overly localizes the Ti-$3d$ and O-2pp states, respectively, driving the system into a non-polar cubic phase even when the small inter-site interaction $V_{Ti,O}$ is added (Fig. S5(f)). A similar trend occurs in eACBN0-based DFT+$U$ calculations, where the combined inclusion of small $U_{Ti}$ and large $U_{O}$ suppresses the polarization and restores cubic symmetry. However, unlike LRT, adding large $V_{Ti,O}$ within the eACBN0 framework recovers the rhombohedral distortion and the associated polarization, consistent with previous findings Choi:2025. Nonetheless, the resulting lattice parameter and band gap remain substantially overestimated.
@@
Figure 8: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZnO without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Zn}$ parameter for 3dd orbitals of Zn ions and inter-site Hubbard $V_{Zn,O}$ parameter between 3dd orbitals of Zn ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap Mang1995ssc, Chen:1998, Reynolds1999prb, Dong2004prb is shown by the dashed blue lines in panel (b).
@@
We first discuss the results for ZnO without the $U_{O}$ correction. As shown in Fig. 8, the LRT method yields an unphysically large self-consistent $U_{Zn}$ exceeding 100 eV. Even without performing the full self-consistent cycle (see Fig. 1), i.e. using the “one-shot” scheme, the LRT estimate already exceeds 25 eV. This well-known limitation of LRT arises from the fully occupied Zn-3dd shell, which leads to spuriously large $U$ values Yu:2014. The “structurally self-consistent” scheme produces a similarly large $U_{Zn}$, as shown in Fig. S6. It is important to emphasize that in ZnO the valence-band maximum is primarily O-2pp in character, while the conduction-band minimum is dominated by Zn-4ss states. As a result, applying $U_{Zn}$ to the Zn-3dd states influences the band gap only indirectly. Indeed, LRT-based DFT+$U$ increases the PBEsol band gap from $\sim 0.7$ eV to 1.94 eV, which remains well below the experimental value of $3.3\sim 3.4$ eV Mang1995ssc, Chen:1998, Reynolds1999prb, Dong2004prb. In contrast, the eACBN0 approach predicts also large but more moderate $U_{Zn}$ of $\sim 16$ eV, resulting in a slightly smaller band gap of 1.52 eV, as shown in Fig. 8(b).
@@ DISCUSSION
We start with the on-site $U_{TM}$ values. For TMOs with partially occupied $3d$ shells, LRT and eACBN0 predict similar $U_{TM}$ values except $U$ for Fe in FeO that differs by 2233 eV. However, when the $3d$ shell is nearly empty or fully occupied, the two methods behave in fundamentally different ways. In particular, for nominally empty $3d$ states, LRT still produces finite and sizable $U_{TM}$ values in the range of 3366 eV, while for fully occupied $3d$ states it yields unphysically large values. The latter is a well-known limitation of the LRT framework Yu:2014: for filled shells, the occupation matrices exhibit vanishing variations under a local electronic perturbation, leading to nearly singular response matrices whose inversion generates spuriously large $U_{TM}$ values. The behavior of eACBN0 is markedly different: It produces $U_{TM}$ values that are finite and in the range of physical values even for closed shells. The root cause of this behavior lies in the definition of the auxiliary objects - the renormalized occupation numbers [Eq. (13)] and renormalized density matrices [Eq. (12)]. These quantities are constructed to ensure that the sum of the Mulliken charge of each atom and the interatomic charge reproduces the correct total electronic population. As a consequence, the Hubbard parameters derived from eACBN0 naturally become very small for nearly empty shells, while for fully filled $3d$ shells they approach values of $\sim 16$ eV. This trend is not mirrored in LRT because of the different physical idea: $U_{TM}$ is computed from the occupation-matrix variations in response to local electronic perturbations. For nearly empty shells, even an infinitesimal local perturbation can induce relatively large fractional changes in the occupation matrix, which translates into sizable $U_{TM}$ values. Conversely, for fully occupied shells, the occupation matrix is extremely rigid, and the small changes captured by the linear-response equations produce nearly vanishing matrix elements. The inversion of these nearly singular matrices amplifies the response artificially, resulting in unphysically large $U_{TM}$ values. Thus, while LRT and eACBN0 can yield comparable results for partially filled shells, their behavior diverges sharply at the two extremes of the $d$-electron occupation spectrum.
@@ METHODS
All calculations were performed using the plane-wave pseudopotential method as implemented in the Quantum ESPRESSO (QE) distribution Giannozzi:2009, Giannozzi:2017, Giannozzi:2020. For the xc functional, we adopted the $\sigma$-GGA with the PBEsol parametrization Perdew:2008, and used pseudopotentials from the GBRV library (v1.5) Garrity:2014. The KS wavefunctions and potentials were expanded in plane waves with kinetic-energy cutoffs of 65 and 780 Ry for the TMOs, and 60 and 600 Ry for ZnO, ZrO2, and BaTiO3. The first BZ was sampled using uniform $\Gamma$-centered $\mathbf{k}$-point meshes: $\times10$$\times$10 for TMOs, $\times12$$\times$7 for ZnO, $\times7$$\times$6 for ZrO2, and $\times8$$\times$8 for BaTiO3. Spin-orbit coupling was neglected. Geometry optimizations were carried out using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm Fletcher:1987, with convergence criteria of $10^{-4}$ Ry for the total energy, $10^{-3}$ Ry/Bohr for forces, and $0.5$ kbar for pressure in TMOs. For ZnO, ZrO2, and BaTiO3, criteria of $10^{-6}$ Ry for the total energy, $10^{-5}$ Ry/Bohr for forces, and $0.5$ kbar for pressure were adopted.
The LRT Hubbard parameters were computed using DFPT Timrov:2018, Timrov:2021 as implemented in the HP code Timrov:2022. Hubbard projectors were constructed from atomic orbitals orthogonalized using the Löwdin scheme Lowdin:1950. Uniform $\Gamma$-centered $\mathbf{k}$- and $\mathbf{q}$-point grids were employed, with sizes of $\times10$$\times$10 and $\times6$$\times$6 for TMOs, $\times12$$\times$7 and $\times6$$\times$3 for ZnO, $\times7$$\times$6 and $\times3$$\times$3 for ZrO2, and $\times8$$\times$8 and $\times4$$\times$4 for BaTiO3, respectively.
The eACBN0 Hubbard parameters were computed using our modified in-house QE package Lee:2020, Yang:2021. For consistency with the LRT approach, the Hubbard projectors were constructed from atomic orbitals orthogonalized using the Löwdin scheme Lowdin:1950. The same $\Gamma$-centered $\mathbf{k}$-point meshes used for the electronic structure calculations were employed, and convergence thresholds of $10^{-10}$ and $10^{-15}$ Ry were applied for the total eACBN0 energy and SCF cycle, respectively.
The $U$ and $V$ parameters were obtained using two self-consistent protocols, summarized in Fig. 1 and Fig. S1 of the SI. In the main text, we report the values computed by iteratively evaluating $U$ and $V$ on the PBEsol-relaxed structures using the LRT and eACBN0 methods until convergence (the “self-consistent” scheme). In the case of ZnO, however, this iterative procedure was limited to two steps due to its tendency to diverge Yu:2014. For completeness, the SI also presents results from the “structurally self-consistent” scheme, in which DFT+UU+$V$ structural optimizations (including Hubbard contributions to forces and stresses Timrov:2020b) are performed, and the Hubbard parameters are recomputed until both the structure and the $U$ and $V$ values are converged.

### v2

@@ RESULTS
Figure 3 extends our analysis by including a Hubbard correction on the oxygen 2pp states ($U_{O}$), computed with both LRT and eACBN0. Within LRT, the resulting $U_{O}$ values are large ($\sim 10$ eV), consistently with previous studies [KirchnerHall:2021]. This is a consequence of the high electronegativity of oxygen: its 2pp shell is nearly fully occupied, which leads to a weak electronic response and therefore a large $U$ in the LRT framework [Gelin:2024, Yu:2014]. Despite these large $U_{O}$ values, the inter-site interaction $V_{TM,O}$ remains small (below 1 eV), indicating that TM–O hybridization has only a minor effect on the electronic structure of these TMOs even when $U_{O}$ is considered. In LRT, $U_{TM}$ and $U_{O}$ appear as diagonal elements, and $V_{TM,O}$ as off-diagonal elements of the matrix obtained by inverting the interacting and non-interacting response matrices, $\chi$ and $\chi_{0}$ [Eqs. (6) and (7)]. Including oxygen among the perturbed subspaces alters the full response matrices, because screening channels involving O are now explicitly taken into account. Once both TM and O sites are perturbed, the structure of the response matrices in DFT+$U$ and DFT+UU+$V$ becomes effectively similar: both schemes capture the same screening pathways involving TM and O states. Consequently, the value of $U_{TM}$ can change when $U_{O}$ is included in the LRT calculation. Comparing Figs. 2 and 3, we observe a slight increase of $U_{TM}$ for CoO and NiO (by 3355%), a substantial increase for FeO (18%), and almost no change for MnO. Conversely, eACBN0 yields smaller $U_{O}$ values than LRT ones, in the range of 232-3 eV. Similar to the case with the $V_{TM,O}$ correction, the converged Hubbard parameters can vary depending on the Hubbard manifolds being considered. In this self-consistent cycle, the parameters from one iteration feed back into the next as a correction term, thereby influencing the KS states and the final converged values. Comparing Figs. 2 and 3 for eACBN0, we find (consistently with LRT) that $U_{TM}$ in CoO increases slightly (by 1122%), while in FeO it slightly decreases (by 1144%) in contrast to LRT. For NiO and MnO, $U_{TM}$ exhibits some variations (small increases, decreases, or no significant change), depending on whether the DFT+$U$ or DFT+UU+$V$ scheme is employed.
@@
Finally, following the protocol in Fig. S1 in the SI [Timrov:2021], we compute the self-consistent Hubbard parameters and their corresponding optimized structural parameters using both LRT and eACBN0. The results are presented in Fig. S3 and S4 and their discussion is in the SI. For both methods, the structural parameters (lattice constants and rhombohedral angles) are generally in good agreement with available experimental data.
@@
For the inter-site interactions between Zr-4dd and O-2pp states ($V_{Zr,O}$), LRT yields values below 1 eV, comparable to those obtained for the TMOs discussed in the previous section. Including these interactions in the DFT+UU+$V$ calculation increases the band gap by 7% relative to DFT+$U$ with only $U_{Zr}$ from LRT. In contrast, eACBN0 produces substantially larger $V_{Zr,O}$ values (2.3–2.5 eV), leading to a 22% band-gap increase compared to DFT+$U$ with only $U_{Zr}$, as shown in Fig. 4(b). Since $V_{Zr,O}$ depends on the interatomic distances between Zr and its nearest-neighbor O atoms, a range of values is reported in Fig. 4(a) rather than a single number. Despite these $V_{Zr,O}$ corrections, both methods still underestimate the experimental band gap by more than 1 eV [Bersch:2008].
@@
Figure 4: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZrO2 without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Zr}$ parameter for 4dd orbitals of Zr ions and inter-site Hubbard $V_{Zr,O}$ parameter between 4dd orbitals of Zr ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap [Bersch:2008] is shown by the dashed blue line in panel (b).
Next, we turn to BaTiO3, which is similar to ZrO2 in the sense that Ti4+ also has a nominal $d^{0}$ configuration. In this case, we consider not only the band gap but also the crystal-structure properties, which are highly sensitive to Hubbard corrections [Gebreyesus:2023, Choi:2025]. BaTiO3 is a well-known ferroelectric material, exhibiting a rhombohedral distortion at low temperatures. However, standard DFT+$U$ with sizable $U$ corrections for Ti-$3d$ states incorrectly stabilizes a centrosymmetric cubic structure, thereby suppressing its intrinsic polarization [Tsunoda2019prm]. However, including the inter-site interactions ($V_{Ti,O}$) between Ti-$3d$ and O-$2p$ states preserves their hybridization, which is essential for maintaining the rhombohedral distortion and thus restoring the polarization [Gebreyesus:2023, Choi:2025].
@@
Figure 6: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for BaTiO3 without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Ti}$ parameter for 3dd orbitals of Ti ions and inter-site Hubbard $V_{Ti,O}$ parameter between 3dd orbitals of Ti ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap [Wemple:1970] is shown by the dashed blue line in panel (b).
@@
When $U_{O}$ is included, the structural behavior changes, as shown in Figs. S5(e)–(h). In this case, LRT fails to capture the ferroelectric nature of BaTiO3: the large $U_{Ti}$ and $U_{O}$ values overly localizes the Ti-$3d$ and O-2pp states, respectively, driving the system into a non-polar cubic phase even when the small inter-site interaction $V_{Ti,O}$ is added (Fig. S5(f)). A similar trend occurs in eACBN0-based DFT+$U$ calculations, where the combined inclusion of small $U_{Ti}$ and large $U_{O}$ suppresses the polarization and restores cubic symmetry. However, unlike LRT, adding large $V_{Ti,O}$ within the eACBN0 framework recovers the rhombohedral distortion and the associated polarization, consistent with previous findings [Choi:2025]. Nonetheless, the resulting lattice parameter and band gap remain substantially overestimated.
@@
Figure 8: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZnO without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Zn}$ parameter for 3dd orbitals of Zn ions and inter-site Hubbard $V_{Zn,O}$ parameter between 3dd orbitals of Zn ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap [Mang1995ssc, Chen:1998, Reynolds1999prb, Dong2004prb] is shown by the dashed blue lines in panel (b).
@@
We first discuss the results for ZnO without the $U_{O}$ correction. As shown in Fig. 8, the LRT method yields an unphysically large self-consistent $U_{Zn}$ exceeding 100 eV. Even without performing the full self-consistent cycle (see Fig. 1), i.e. using the “one-shot” scheme, the LRT estimate already exceeds 25 eV. This well-known limitation of LRT arises from the fully occupied Zn-3dd shell, which leads to spuriously large $U$ values [Yu:2014]. The “structurally self-consistent” scheme produces a similarly large $U_{Zn}$, as shown in Fig. S6. It is important to emphasize that in ZnO the valence-band maximum is primarily O-2pp in character, while the conduction-band minimum is dominated by Zn-4ss states. As a result, applying $U_{Zn}$ to the Zn-3dd states influences the band gap only indirectly. Indeed, LRT-based DFT+$U$ increases the PBEsol band gap from $\sim 0.7$ eV to 1.94 eV, which remains well below the experimental value of $3.3\sim 3.4$ eV [Mang1995ssc, Chen:1998, Reynolds1999prb, Dong2004prb]. In contrast, the eACBN0 approach predicts also large but more moderate $U_{Zn}$ of $\sim 16$ eV, resulting in a slightly smaller band gap of 1.52 eV, as shown in Fig. 8(b).
@@ DISCUSSION
We start with the on-site $U_{TM}$ values. For TMOs with partially occupied $3d$ shells, LRT and eACBN0 predict similar $U_{TM}$ values except $U$ for Fe in FeO that differs by 2233 eV. However, when the $3d$ shell is nearly empty or fully occupied, the two methods behave in fundamentally different ways. In particular, for nominally empty $3d$ states, LRT still produces finite and sizable $U_{TM}$ values in the range of 3366 eV, while for fully occupied $3d$ states it yields unphysically large values. The latter is a well-known limitation of the LRT framework [Yu:2014]: for filled shells, the occupation matrices exhibit vanishing variations under a local electronic perturbation, leading to nearly singular response matrices whose inversion generates spuriously large $U_{TM}$ values. The behavior of eACBN0 is markedly different: It produces $U_{TM}$ values that are finite and in the range of physical values even for closed shells. The root cause of this behavior lies in the definition of the auxiliary objects - the renormalized occupation numbers [Eq. (13)] and renormalized density matrices [Eq. (12)]. These quantities are constructed to ensure that the sum of the Mulliken charge of each atom and the interatomic charge reproduces the correct total electronic population. As a consequence, the Hubbard parameters derived from eACBN0 naturally become very small for nearly empty shells, while for fully filled $3d$ shells they approach values of $\sim 16$ eV. This trend is not mirrored in LRT because of the different physical idea: $U_{TM}$ is computed from the occupation-matrix variations in response to local electronic perturbations. For nearly empty shells, even an infinitesimal local perturbation can induce relatively large fractional changes in the occupation matrix, which translates into sizable $U_{TM}$ values. Conversely, for fully occupied shells, the occupation matrix is extremely rigid, and the small changes captured by the linear-response equations produce nearly vanishing matrix elements. The inversion of these nearly singular matrices amplifies the response artificially, resulting in unphysically large $U_{TM}$ values. Thus, while LRT and eACBN0 can yield comparable results for partially filled shells, their behavior diverges sharply at the two extremes of the $d$-electron occupation spectrum.
@@ METHODS
All calculations were performed using the plane-wave pseudopotential method as implemented in the Quantum ESPRESSO (QE) distribution [Giannozzi:2009, Giannozzi:2017, Giannozzi:2020]. For the xc functional, we adopted the $\sigma$-GGA with the PBEsol parametrization [Perdew:2008], and used pseudopotentials from the GBRV library (v1.5) [Garrity:2014]. The KS wavefunctions and potentials were expanded in plane waves with kinetic-energy cutoffs of 65 and 780 Ry for the TMOs, and 60 and 600 Ry for ZnO, ZrO2, and BaTiO3. The first BZ was sampled using uniform $\Gamma$-centered $\mathbf{k}$-point meshes: $\times10$$\times$10 for TMOs, $\times12$$\times$7 for ZnO, $\times7$$\times$6 for ZrO2, and $\times8$$\times$8 for BaTiO3. Spin-orbit coupling was neglected. Geometry optimizations were carried out using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [Fletcher:1987], with convergence criteria of $10^{-4}$ Ry for the total energy, $10^{-3}$ Ry/Bohr for forces, and $0.5$ kbar for pressure in TMOs. For ZnO, ZrO2, and BaTiO3, criteria of $10^{-6}$ Ry for the total energy, $10^{-5}$ Ry/Bohr for forces, and $0.5$ kbar for pressure were adopted.
The LRT Hubbard parameters were computed using DFPT [Timrov:2018, Timrov:2021] as implemented in the HP code [Timrov:2022]. Hubbard projectors were constructed from atomic orbitals orthogonalized using the Löwdin scheme [Lowdin:1950]. Uniform $\Gamma$-centered $\mathbf{k}$- and $\mathbf{q}$-point grids were employed, with sizes of $\times10$$\times$10 and $\times6$$\times$6 for TMOs, $\times12$$\times$7 and $\times6$$\times$3 for ZnO, $\times7$$\times$6 and $\times3$$\times$3 for ZrO2, and $\times8$$\times$8 and $\times4$$\times$4 for BaTiO3, respectively.
The eACBN0 Hubbard parameters were computed using our modified in-house QE package [Lee:2020, Yang:2021]. For consistency with the LRT approach, the Hubbard projectors were constructed from atomic orbitals orthogonalized using the Löwdin scheme [Lowdin:1950]. The same $\Gamma$-centered $\mathbf{k}$-point meshes used for the electronic structure calculations were employed, and convergence thresholds of $10^{-10}$ and $10^{-15}$ Ry were applied for the total eACBN0 energy and SCF cycle, respectively.
The $U$ and $V$ parameters were obtained using two self-consistent protocols, summarized in Fig. 1 and Fig. S1 of the SI. In the main text, we report the values computed by iteratively evaluating $U$ and $V$ on the PBEsol-relaxed structures using the LRT and eACBN0 methods until convergence (the “self-consistent” scheme). In the case of ZnO, however, this iterative procedure was limited to two steps due to its tendency to diverge [Yu:2014]. For completeness, the SI also presents results from the “structurally self-consistent” scheme, in which DFT+UU+$V$ structural optimizations (including Hubbard contributions to forces and stresses [Timrov:2020b]) are performed, and the Hubbard parameters are recomputed until both the structure and the $U$ and $V$ values are converged.